# LongLLaDA: Unlocking Long Context Capabilities in Diffusion LLMs

### v1

@@ Abstract
Large Language Diffusion Models, or diffusion LLMs, have emerged as a significant focus in NLP research, with substantial effort directed toward understanding their scalability and downstream task performance. However, their long-context capabilities remain unexplored, lacking systematic analysis or methods for context extension. In this work, we present the first systematic investigation comparing the long-context performance of diffusion LLMs and traditional auto-regressive LLMs. We first identify a unique characteristic of diffusion LLMs, unlike auto-regressive LLMs, they maintain remarkably stable perplexity during direct context extrapolation. Furthermore, where auto-regressive models fail outright during the Needle-In-A-Haystack task with context exceeding their pretrained length, we discover diffusion LLMs exhibit a distinct “local perception” phenomenon, enabling successful retrieval from recent context segments. We explain both phenomena through the lens of Rotary Position Embedding (RoPE) scaling theory. Building on these observations, we propose LongLLaDA, a training-free method that integrates LLaDA with the NTK-based RoPE extrapolation. Our results validate that established extrapolation scaling laws remain effective for extending the context windows of diffusion LLMs. Furthermore, we identify long-context tasks where diffusion LLMs outperform auto-regressive LLMs and others where they fall short. Consequently, this study establishes the first context extrapolation method for diffusion LLMs while providing essential theoretical insights and empirical benchmarks critical for advancing future research on long-context diffusion LLMs.
@@ 1 Introduction
(a) LLaDA-8B-Base
(b) LLaMA3-8B-Base
@@
Recently, diffusion LLMs have become a widely discussed topic in Natural Language Processing research (Nie et al., 2025; Ye et al., 2025). They are regarded as a potential solution to key limitations of traditional auto-regressive LLMs, including the reversal curse (Berglund et al., 2023), complex reasoning (Dziri et al., 2023), long-term planning, and maintaining coherence across extended contexts (Bachmann & Nagarajan, 2024; Ye et al., 2024; 2025). Significant research efforts have focused on validating their scalability (Nie et al., 2025; Ye et al., 2025), adapting them for multimodality (Yang et al., 2025; You et al., 2025; Yu et al., 2025), applying them to reasoning tasks (Zhao et al., 2025; Huang et al., 2025; Zhu et al., 2025), and optimizing their efficiency (Ma et al., 2025; Hu et al., 2025; Wu et al., 2025). However, the long-context capabilities of diffusion LLMs, specifically their performance and potential for length extrapolation, remain unexplored.
@@ 6 Related Work
Recently, Large Language Diffusion Models, or diffusion LLMs, have become a widely discussed topic in NLP research. After the theoretical simplification (Sahoo et al., 2024; Ou et al., 2024) and fine-tuning verification (Gong et al., 2024), researchers scale the size of diffusion LLMs to billions of parameters (Nie et al., 2024; 2025; Ye et al., 2025) and demonstrate that diffusion LLMs can achieve comparable results with more promising performance in the reversal curse (Berglund et al., 2023). These immediately attract the attention of many more researchers. Significant research efforts have focused on adapting diffusion LLMs for multimodality, such as MMaDA (Yang et al., 2025), LLaDA-V (You et al., 2025), and LaViDa (Li et al., 2025), applying them to reasoning tasks, such as d1 (Zhao et al., 2025), DCoLTHuang et al. (2025), and LLaDA-1.5 (Zhu et al., 2025), and optimizing their efficiency (Ma et al., 2025; Hu et al., 2025; Wu et al., 2025), including dKV-Cache (Ma et al., 2025), Dimple (Yu et al., 2025), dLLM-Cache (Liu et al., ), FreeCache (Hu et al., 2025), Fast-dLLM (Wu et al., 2025), and so on. However, there is still no discussion on the long-context capability of diffusion LLMs.
@@
Length extrapolation, or length generalization, or context extension, is an important issue for LLMs (Press et al., 2022). The mainstream extrapolation research mainly focuses on adjusting position embedding, especially the widely used RoPE (Su et al., 2021). For example, Linear PI (Chen et al., 2023) first achieves LLMs’ length extrapolation by scaling position indices to the pre-training range with little fine-tuning. The NTK method (bloc97, 2023b; a; Peng et al., 2023) then scales the rotary base in RoPE (Su et al., 2021) to achieve plug-and-play length extrapolation. Subsequently, amplifying the rotary base and training on longer lengths has become the dominant approach for length extrapolation (Rozière et al., 2023; Xiong et al., 2023; Liu et al., 2023b; Ding et al., 2024). In addition, ReRoPE (Su, 2023), ReAttention (Liu et al., 2024b), and DCA (An et al., 2024a; b) also achieve plug-and-play extrapolation by limiting the relative position. In this paper, we still focus on the length extrapolation via NTK scaling (bloc97, 2023b; Liu et al., 2023b) in the inference stage, and try to reveal and explain the similarities and differences in length extrapolation between diffusion-based and auto-regressive LLM.

### v2

@@ Abstract
Large Language Diffusion Models, or diffusion LLMs, have emerged as a significant focus in NLP research, with substantial effort directed toward understanding their scalability and downstream task performance. However, their long-context capabilities remain unexplored, lacking systematic analysis or methods for context extension. In this work, we present the first systematic investigation comparing the long-context performance of diffusion LLMs and traditional auto-regressive LLMs. We first identify a unique characteristic of diffusion LLMs, unlike auto-regressive LLMs, they maintain remarkably stable perplexity during direct context extrapolation. Moreover, where auto-regressive models fail outright during the Needle-In-A-Haystack task with context exceeding their pretrained length, we discover diffusion LLMs exhibit a distinct “local perception” phenomenon, enabling successful retrieval from recent context segments. We explain both phenomena through the lens of Rotary Position Embedding (RoPE) scaling theory. Building on these observations, we propose LongLLaDA, a training-free method that integrates LLaDA with the NTK-based RoPE extrapolation. Our results validate that established extrapolation scaling laws remain effective for extending the context windows of diffusion LLMs. Furthermore, we identify long-context tasks where diffusion LLMs outperform auto-regressive LLMs and others where they fall short. Consequently, this study establishes the first length extrapolation method for diffusion LLMs while providing essential theoretical insights and empirical benchmarks critical for advancing future research on long-context diffusion LLMs. The code is available at https://github.com/OpenMOSS/LongLLaDA.
@@ 1 Introduction
(a) Comparison of Perplexity
(b) Comparison of Retrieval Accuracy
@@
Recently, diffusion LLMs have become widely discussed in Natural Language Processing research (Nie et al., 2025; Ye et al., 2025). They are regarded as a potential solution to key limitations of traditional auto-regressive LLMs (Touvron et al., 2023; Sun et al., 2024), including the reversal curse (Berglund et al., 2023), complex reasoning (Dziri et al., 2023), and maintaining coherence across extended contexts (Bachmann & Nagarajan, 2024; Ye et al., 2024; 2025). Significant research efforts have focused on validating their scalability (Nie et al., 2025; Ye et al., 2025), adapting them for multimodality (Yang et al., 2025; You et al., 2025; Yu et al., 2025), applying them to reasoning tasks (Zhao et al., 2025; Huang et al., 2025; Zhu et al., 2025), and optimizing their efficiency (Ma et al., 2025; Hu et al., 2025; Wu et al., 2025). However, the long-context capabilities of diffusion LLMs, specifically their performance and potential for length extrapolation, remain unexplored.
@@ 6 Related Work
Recently, Large Language Diffusion Models, or diffusion LLMs, have become a widely discussed topic in NLP research. After the theoretical simplification (Sahoo et al., 2024; Ou et al., 2024; Shi et al., 2024) and empirical verification (He et al., 2022; Gong et al., 2024), researchers scale the size of diffusion LLMs to billions of parameters (Nie et al., 2024; 2025; Ye et al., 2025) and demonstrate that diffusion LLMs can achieve comparable results with more promising performance in the reversal curse (Berglund et al., 2023). These immediately attract the attention of many more researchers. Significant research efforts have focused on adapting diffusion LLMs for multimodality, such as MMaDA (Yang et al., 2025), LLaDA-V (You et al., 2025), and LaViDa (Li et al., 2025), applying them to reasoning tasks, such as d1 (Zhao et al., 2025), DCoLTHuang et al. (2025), and LLaDA-1.5 (Zhu et al., 2025), and optimizing their efficiency (Ma et al., 2025; Hu et al., 2025; Wu et al., 2025), including dKV-Cache (Ma et al., 2025), Dimple (Yu et al., 2025), dLLM-Cache (Liu et al., ), FreeCache (Hu et al., 2025), Fast-dLLM (Wu et al., 2025), and so on. However, there is still no discussion on the long-context capability of diffusion LLMs.
@@
Length extrapolation, or length generalization, or context extension, is an important issue for LLMs (Press et al., 2022; Liu et al., 2025). The mainstream extrapolation research mainly focuses on adjusting position embedding, especially the widely used RoPE (Su et al., 2021). For example, Linear PI (Chen et al., 2023) first achieves LLMs’ length extrapolation by scaling position indices to the pre-training range with little fine-tuning. The NTK method (bloc97, 2023b; a; Peng et al., 2023) then scales the rotary base in RoPE (Su et al., 2021) to achieve plug-and-play length extrapolation. Subsequently, amplifying the rotary base and training on longer lengths has become the dominant approach for length extrapolation (Rozière et al., 2023; Xiong et al., 2023; Liu et al., 2023b; Ding et al., 2024). In addition, ReRoPE (Su, 2023), ReAttention (Liu et al., 2024b), and DCA (An et al., 2024a; b) also achieve plug-and-play extrapolation by limiting the relative position. In this paper, we still focus on the length extrapolation via NTK scaling (bloc97, 2023b; Liu et al., 2023b) in the inference stage, and try to reveal and explain the similarities and differences in length extrapolation between diffusion-based and auto-regressive LLM.